OpenStudy (anonymous):

Richard is conducting an experiment where he flips a two-sided coin 3 times. How many different combinations of outcomes could the experiment have? A.16 B.8 C.4 D.10

OpenStudy (bibby):

there are 2 possibilities each time \(2*2*2=2^3=8\)

OpenStudy (anonymous):

yeah I got 8 XD

OpenStudy (anonymous):

o-o

OpenStudy (bibby):

o~o

OpenStudy (anonymous):

ITS A BANANANANANANA O-O

OpenStudy (anonymous):

0-0

OpenStudy (anonymous):

dies instantly

OpenStudy (bibby):

rip, I'm gonna head off to another question peace brotha

OpenStudy (anonymous):

o-O

OpenStudy (anonymous):

peace

OpenStudy (anonymous):

sister >.>

OpenStudy (mathstudent55):

Here's a tree diagram that shows the 8 possible different outcomes.